OpenStudy (anonymous):

The following system is solved by substitution. Which expression is substituted for z in the second and third equations? x + 5y – z = 6 3x – y – 2z = 4 3x – 2y + 3z = 13

OpenStudy (anonymous):

Well, x+5y-z = 6 so it follows: x-x+5y-5y-z = 6-x-5y (x and 5y cancel) -z = 6-x-5y (divide by -1 both sides) and then: z = -6+x+5y is what you use for the other two equations for z.